I'm reading that post, and while fascinating, I don't understand the entropy calculation.
The digit chance is easy:
For a single digit to appear one needs 1/3 * 1/10 = 1/30.
For a single letter to appear one needs 1/3 * 1/26 = 1/78.
But the bits of entropy throw me off. 26 + 26 + 10 = 62, which is 2^5.954. But that is for a uniform distribution. The writer states that it actually is 2^5.826, or 1/~56.7. I don't get how they to that number.